The percentage y (of total personal consumption) an individual spends on food is approximately y = 33 x^-0.63 percentage points ( 2.5 ≤ x ≤ 4.5 ) where x is the percentage the individual spends on recreation.? A college student finds that he is spending x = 3.5 + 0.1 t percent of his personal consumption on recreation, where t is time in months since January 1. Use direct substitution to express the percentage y as a function of time t (do not simplify the expression). (NOTE: January 1 is represented by t = 0.) y(t) = Use the chain rule to estimate how fast the percentage he spends on food is changing on February 1. (Round your answer to two decimal places.)
Answer #1

-0.63 -1.63 y(t) = (x64),-0.62 = (3.5 +01+) y(t) = -0.63 (3.5 +0.lt) at tur tie February 1 yet = -0.63 3.67 = -0.1192 te 7. h